OpenStudy (anonymous):

9. What digit appears in the units place in the number obtained when 2^320 is multiplied out? A. 0 B. 2 C. 4 D. 6 E. 8

OpenStudy (anonymous):

If we wanted the last digit of 2 999 , note that 996 is a multiple of 4 . So at 996 we get a 6 . Now count forward: 2 , 4 , 8 : the answer is 8 . Or else 1000 is a multiple of 4 . Go backwards one step from 6 : the last digit is 8 .

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

well im still loading the concept ....

OpenStudy (unklerhaukus):

\[2^1=2\]\[2^2=4\]\[2^3=8\]\[2^4={16}\]\[2^5=32\]\[2^6=64\]\[2^7=128\]\[2^8=256\]\[\dots\] we notice last digit cycles, 2,4,8,6,2,4,8,6,2,4,6... the pattern repeats every 4 terms

OpenStudy (anonymous):

oh ok i c that ...

OpenStudy (anonymous):

right that's helpful

OpenStudy (unklerhaukus):

so the last digit will be the same for \[2^{320},2^{316},...\]

OpenStudy (anonymous):

uh huh ...

OpenStudy (unklerhaukus):

\(320\) is 4 times 79 + 4 \(2^{320}\) has the same last digit as \(2^4\)

OpenStudy (anonymous):

right ....thanx alot